OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

you need the slope of \[4x-3y=-12\] the slope of \[ax+by=c\] is \[m=-\frac{a}{b}\]

OpenStudy (anonymous):

okso the slope is 4/3 @misty1212

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

now \[y-4=\frac{4}{3}(x+3)\] solve for \(y\)

OpenStudy (anonymous):

ok y-4+4 =4/3x +4+4

OpenStudy (anonymous):

the answer is a @misty1212 right

OpenStudy (misty1212):

it is correct when you write \[y=\frac{4}{3}x+8\]

OpenStudy (anonymous):

so that is the answer a???

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

lol yeah